# End to End Collaborative Synthetic Data Generation

### v1

@@ 2 Preliminaries
Secure Multi-Party Computation (MPC). MPC refers to a class of cryptographic approaches that enable multiple distrustful parties to privately compute a function without revealing their private inputs to any other entity [20, 21]. We consider secret sharing based MPC protocols. Our framework employs the “MPC as a service” paradigm, where the data custodians delegate their private computations to a set of non-colluding servers, which we refer to as MPC servers. Each data custodian splits their private input $v$ into secret shares denoted as $[\![v]\!]$, and sends the shares to the MPC servers, following the given MPC scheme. Though $v$ can be reconstructed when the shares are combined, none of the MPC servers can individually learn the value of $v$. The MPC servers then proceed to jointly perform computations over the secret-shared data such as preprocessing the data, training a machine learning model, evaluating models and generating synthetic data. Since all computations are performed on the secret shares, the servers cannot infer the input values nor intermediate results, thus ensuring that MPC guarantees input privacy. We refer to Appendix A for details on threat models and on how to perform addition and multiplication with a replicated secret sharing scheme, as well as an overview of MPC primitives that we use in this paper.
[…]
@@ 3 Secure Generation and Publishing of Synthetic Data
2:Data custodians, $\mathbf{C}$, secret share their respective datasets with MPC servers, $\mathbf{S}$, as $[\![D_{i}]\!]$ and thresholds of data quality $[\![T_{i}]\!]$
3:$\mathbf{S}$ run setup
4:$[\![D]\!]\leftarrow\pi_{\mathsf{CONCAT}}([\![D_{i}]\!]\,|\,\forall i\in|\mathbf{S}|)$
@@
13:         $[\![D_{train}]\!],[\![D_{test}]\!]\leftarrow$ getData($[\![D]\!]$)
14:         $[\![D_{train}]\!]\leftarrow\pi_{\mathsf{PREPROCESS}}([\![D_{train}]\!],(\epsilon_{p},\delta_{p}))$
15:         $[\![\hat{D}_{train}]\!]\leftarrow\pi_{\mathsf{SDG}}([\![D_{train}]\!],(\epsilon_{s},\delta_{s}),\mathbf{H_{k}})$
16:         $[\![r]\!]\leftarrow\pi_{\mathsf{EVAL}}([\![\hat{D}_{train}]\!],[\![D_{test}]\!],[\![D_{train}]\!])$
17:         $[\![R[k]]\!]\leftarrow[\![r]\!]$
@@
19:     $[\![M]\!]\leftarrow\pi_{\mathsf{AVG}}([\![R[k]]\!]|\forall k)$
@@
21:     for every data custodian $c$ get threshold $[\![T_{c}]\!]$ do
22:         votes $\leftarrow$ votes $-([\![M]\!<[\![T_{c}]\!]))$
@@
28:     $[\![{D}]\!]\leftarrow\pi_{\mathsf{PREPROCESS}}([\![D]\!],(\epsilon_{p},\delta_{p}))$
29:     $[\![\hat{D}]\!]\leftarrow\pi_{\mathsf{SDG}}([\![D]\!],(\epsilon_{s},\delta_{s}),\mathbf{H_{s}})$
30:     $[\![\hat{D}]\!]\leftarrow\pi_{\mathsf{REV-PREPROCESS}}([\![\hat{D}]\!])$
@@
The data custodians $\mathbf{C}$ secret share their private data $[\![D_{i}]\!]$ and the desired quality of synthetic data as thresholds $[\![T_{i}]\!]$ to MPC servers $\mathbf{S}$ on Line 1, while servers set up based on the secret-shared inputs received from the data custodians and public inputs such as the dimensions of the data and the set of parameters to tune for. $\mathbf{S}$ then run $\pi_{\mathsf{CONCAT}}$ on Line 3 to concat the secret shares of the data from all $\mathbf{C}$ and hold the secret shares of $[\![D]\!]$ at the end of this protocol, hence, mimicking the centralized setup, as if all the data was in one location.
@@
Parameter tuning is performed iteratively from Lines 6 – 25 until one of the following conditions is met: the hyperparameter set is exhausted, the maximum number of iterations is reached, or the desired quality of synthetic data is achieved. [^7] This can be easily modified to exhaustively search for all parameters, and then choose the best hyperparameter for which we obtained the best evaluation metrics. This would require large number of secure comparisons - comparing with thresholds and comparing with all the sets of averaged evaluation metrics. Each iteration begins with K-fold (Lines 8 – 17), creating train and test splits based on indices only with getData() on Line 12. $\mathbf{S}$ then run MPC protocol $\pi_{\mathsf{PREPROCESS}}$ to preprocess the secret shares of training data on Line 13 followed by an MPC protocol for $\pi_{\mathsf{SDG}}$ to train an SDG model on secret shares of preprocessed training data with the set of hyperparameter $\mathbf{H_{k}}$. On line 15, $\mathbf{S}$ run $\pi_{\mathsf{EVAL}}$ to compute secret shares of the pre-defined evaluation metrics $[\![r]\!]$ based on secret shares of both the real $[\![D_{train}]\!],[\!D_{test}]\!]$ and generated synthetic data $[\![\hat{D}_{train}]\!]$. The value of $[\![r]\!]$ is never published and hence need not be DP. Note that all of these steps perform computation on combined real data and hence are done in MPC. On line 18, $\mathbf{S}$ compute secret shares of K-fold evaluation $[\![M]\!]$ using $\pi_{\mathsf{AVG}}$.
@@
Lines 19 – 23 determine if $[\![M]\!]$ meets the thresholds of all the data custodians. To do so, assume that $[\![M]\!]$ meets the threshold for all $\mathbf{C}$ on Line 19, i.e., votes $=|\mathbf{C}|$. Line 21 decrements the votes if it does not meet the threshold of a data custodian, finally checking if the number of votes still equals $|\mathbf{C}|$ on Line 23. If yes, it is set to publish on Line 23, selecting the hyperparameters for this iteration as the final set of hyperparamters $\mathbf{H_{s}}$. The voting is done in a privacy-preserving way – (a) it does not require release of any evaluation metrics; and (b) the thresholds of each data custodian and the number of votes are kept confidential. Finally, on Line 27 – 29, the SDG pipeline is run on the combined data $[\![D]\!]$, including reverse preprocessing on Line 29 to match the format of the real data that was secret shared initially. Line 30 publishes generated synthetic data of the same length as the combined data, which can be changed in the framework (e.g. on Line 28). Note that the framework is modular in nature and can be switched to suit any specific preprocessing or SDG algorithm or evaluation metrics [^8] The code for our framework is available at https://https://anonymous.4open.science/endsdg. Note that it is work in progress.. Once we have the MPC protocols in-place, one can easily modify to change the method of parameter search as long as it does not require access to private data.
@@ 4.1 Secure Preprocessing
1:Secret shares of one gene feature column of $[\![D^{g}]\!]$ for gene $g$, the number of samples $N$
2:Secret shares of one gene feature binned column $[\![D^{g}]\!]$ for gene $g$, secret shares of mean $[\![m^{g}\!]$ for gene $g$ for inverse discretization done later
@@
4:Create a copy of $[\![D^{g}]\!]$ as $[\![{D^{g}_{copy}}]\!]$; Initialize $[\![Q]\!]$ of length 3
5:$[\![\widetilde{D^{g}}]\!]\leftarrow\pi_{\mathsf{SORT}}([\![D^{g}]\!])$triangleright MPC protocol for sorting based on radix sort; see MP-SPDZ [31]
@@
8:     $[\![Q[$$r$$]]\!]\leftarrow[\![\widetilde{D^{g}}[\lfloor$pos]]$\rfloor$]]\!]⌋ ] ] ] + $f\cdot$($[\![\widetilde{D^{g}}[\lfloor$pos+1]]$\rfloor$+1]]\!]⌋ + 1 ] ] ]-$[\![\widetilde{D^{g}}[\lfloor$pos]]$\rfloor$]]\!]⌋ ] ] ])
@@
10:for $i\leftarrow 1$ to $N$ do $\triangleright$ Binning data into 4 bins with values 0,1,2,3
[…]
12:     $[\![D^{g}[i]]\!]\leftarrow 3-[\![c_{0}]\!]-[\![c_{1}]\!]-[\![c_{2}]\!]$
@@
15:for $i\leftarrow 1$ to $N$ do $\triangleright$ Computing mean for inverse discretization
16:     for $b\leftarrow 0$ to $3$ do
17:         $[\![c]\!]\leftarrow\pi_{\mathsf{EQ}}([\![D^{g}[i]]\!],b)$
18:         $[\![m^{g}[b]\!]\leftarrow[\![m^{g}[b]\!]+\pi_{\mathsf{MUL}}([\![c]\!],[\![D^{g}_{copy}[i]]\!]$)
[…]
@@
[…]
Description of $\pi_{\mathsf{INV-BIN}}$. $\pi_{\mathsf{INV-BIN}}$ takes as input secret shares of binned $[\![{D}]\!]$ and secret shares of all mean values $[\![m^{g}]\!]$ and outputs the de-binned data. We replace the value $g_{i}$ with the mean following $\sum_{i=0}^{3}m^{g}[b]\cdot(g_{i}==b)$ on Lines 4–5. We note that line 4 can be further optimized by using the indicator polynomials described in Section 4.2.
@@
1:Secret shares of binned data of $[\![D]\!]$, the number of samples $N$, secret shares of mean $[\![m^{g}\!]$ for all genes
2:Secret shares of de-binned data $[\![D]\!]$
3:for $i\leftarrow 1$ to $N$ do $\triangleright$ Binned Data into 4 bins with values 0,1,2,3
@@
5:         $[\![x]\!]\leftarrow[\![D[i][g]]\!]$
6:         $[\![c_{0}]\!]\leftarrow\pi_{\mathsf{EQ}}([\![x]\!],0)$; $[\![c_{1}]\!]\leftarrow\pi_{\mathsf{EQ}}([\![x]\!],1)$; $[\![c_{2}]\!]\leftarrow\pi_{\mathsf{EQ}}([\![x]\!],2)$; $[\![c_{3}]\!]\leftarrow\pi_{\mathsf{EQ}}([\![x]\!],3)$
7:         $[\![D[i][g]]\!]\leftarrow\pi_{\mathsf{MUL}}([\![c_{0}]\!],[\![m^{g}\!][0])+\pi_{\mathsf{MUL}}([\![c_{1}]\!],[\![m^{g}\!][1])+\pi_{\mathsf{MUL}}([\![c_{2}]\!],[\![m^{g}\!][2])+\pi_{\mathsf{MUL}}([\![c_{3}]\!],[\![m^{g}\!][3])$
@@ 4.2 Secure SDG
$\begin{array}[]{lcl}I_{0}(x)=\frac{(1-x)(2-x)(3-x)}{6}&\phantom{xxx}&I_{1}(x)=\frac{x(2-x)(3-x)}{2}\\ &&\\ I_{2}(x)=\frac{x(x-1)(3-x)}{2}&&I_{3}(x)=\frac{x(x-1)(x-2)}{6}\\ \end{array}$
@@
1:Secret shares of binned data $[\![D]\!]$, scale $\sigma$, number of patients $N$, Domain of gene features $\Omega_{g}={0,1,2,3}$ and classes $\Omega_{y}={0,1,2,3,4}$
2:Secret shares of noisy marginals $[\![\mu]\!]$ for 1-way and selected 2-way marginals
3:Initialize matrix $[\![\mu_{g}]\!]$ of size $[d,4]$, array $[\![\mu_{y}]\!]$ of size $[5]$ and matrix $[\![\mu_{g,y}]\!]$ of size $[d,20]$
4:for $i\leftarrow 1$ to $N$ do
5:     $[\![x]\!]\leftarrow[\![D[i][y]]\!]$triangleright Compute 1-way marginal for the label
6:     Initialize array $[\![L]\!]$ of size 5
7:     $[\![s_{0}]\!]\leftarrow[\![x]\!]$; $[\![s_{1}]\!]\leftarrow[\![x]\!]-1$; $[\![s_{2}]\!]\leftarrow[\![x]\!]-2$; $[\![s_{3}]\!]\leftarrow[\![x]\!]-3$; $[\![s_{4}]\!]\leftarrow[\![x]\!]-4$;
8:     $[\![L[0]]\!]\leftarrow\pi_{\mathsf{MUL}}([\![s_{1}]\!],[\![s_{2}]\!],[\![s_{3}]\!],[\![s_{4}]\!],(1/24))$
9:     $[\![L[1]]\!]\leftarrow\pi_{\mathsf{MUL}}([\![s_{0}]\!],[\![s_{1}]\!],[\![s_{3}]\!],[\![s_{4}]\!],(1/6))$
10:     $[\![L[2]]\!]\leftarrow\pi_{\mathsf{MUL}}([\![s_{0}]\!],[\![s_{1}]\!],[\![s_{2}]\!],[\![s_{4}]\!],(1/2))$
11:     $[\![L[3]]\!]\leftarrow\pi_{\mathsf{MUL}}([\![s_{0}]\!],[\![s_{1}]\!],[\![s_{2}]\!],[\![s_{4}]\!],(1/6))$
12:     $[\![L[4]]\!]\leftarrow\pi_{\mathsf{MUL}}([\![s_{0}]\!],[\![s_{1}]\!],[\![s_{2}]\!],[\![s_{3}]\!],(1/24))$
@@
17:         $[\![x]\!]\leftarrow[\![D[i][g]]\!]$
18:         Initialize array $[\![G]\!]$ of size 5
19:         $[\![s_{1}]\!]\leftarrow 1-[\![x]\!]$; $[\![s_{11}]\!]\leftarrow[\![x]\!]-1$; $[\![s_{2}]\!]\leftarrow 2-[\![x]\!]$; $[\![s_{21}]\!]\leftarrow[\![x]\!]-2$; $[\![s_{3}]\!]\leftarrow 3-[\![x]\!]$; $[\![s_{31}]\!]\leftarrow[\![x]\!]-3$;
20:         $[\![G[0]]\!]\leftarrow\pi_{\mathsf{MUL}}([\![s_{1}]\!],[\![s_{2}]\!],[\![s_{3}]\!],(1/6))$
21:         $[\![G[1]]\!]\leftarrow\pi_{\mathsf{MUL}}([\![x]\!],[\![s_{2}]\!],[\![s_{3}]\!],(1/2))$
22:         $[\![G[2]]\!]\leftarrow\pi_{\mathsf{MUL}}([\![x]\!],[\![s_{11}]\!],[\![s_{3}]\!],(1/2))$
23:         $[\![G[3]]\!]\leftarrow\pi_{\mathsf{MUL}}([\![x]\!],[\![s_{11}]\!],[\![s_{22}]\!],(1/6))$
@@
35:for every value $[\![x]\!]$ in $[\![\mu_{g}]\!]$, $[\![\mu_{l}]\!]$, $[\![\mu_{l}g]\!]$ do $\triangleright$ Adding noise to guarantee DP
36:     $[\![x]\!]\leftarrow[\![x]\!]+\sigma\cdot\pi_{\mathsf{GAUSS}}(0,1)$
@@
$\begin{array}[]{lcl}I^{\prime}_{0}(y)=\frac{(y-1)(y-2)(y-3)(y-4)}{24}&\phantom{xxx}&I^{\prime}_{1}(y)=\frac{y(y-2)(y-3)(y-4)}{6}\\ &&\\ I^{\prime}_{2}(y)=\frac{y(y-1)(y-3)(y-4)}{2}&&I^{\prime}_{3}(y)=\frac{y(y-1)(y-2)(y-4)}{6}\\ &&\\ I^{\prime}_{4}(y)=\frac{y(y-1)(y-2)(y-3)}{24}&&\\ \end{array}$

### v2

@@ 2 Preliminaries
Secure Multi-Party Computation (MPC). MPC refers to a class of cryptographic approaches that enable multiple distrustful parties to privately compute a function without revealing their private inputs to any other entity [20, 21]. We consider secret sharing based MPC protocols. Our framework employs the “MPC as a service” paradigm, where the data custodians delegate their private computations to a set of non-colluding servers, which we refer to as MPC servers. Each data custodian splits their private input $v$ into secret shares denoted as [[v]][\![v]\!], and sends the shares to the MPC servers, following the given MPC scheme. Though $v$ can be reconstructed when the shares are combined, none of the MPC servers can individually learn the value of $v$. The MPC servers then proceed to jointly perform computations over the secret-shared data such as preprocessing the data, training a machine learning model, evaluating models and generating synthetic data. Since all computations are performed on the secret shares, the servers cannot infer the input values nor intermediate results, thus ensuring that MPC guarantees input privacy. We refer to Appendix A for details on threat models and on how to perform addition and multiplication with a replicated secret sharing scheme, as well as an overview of MPC primitives that we use in this paper.
Differential Privacy (DP). Consider two neighboring datasets $D$ and $D^{\prime}$ that differ by a single record. [^6] DD can be obtained by ether adding or removing a record in $D^{\prime}$. A randomized algorithm $F$ is said to be $(\epsilon,\delta)$-DP if, for any pair of neighboring datasets $D$ and $D^{\prime}$ and for all subsets $O$ in the range of the output of $F$, it holds that $\mbox{Pr}(F(D)\in O)\leq e^{\epsilon}\cdot\mbox{Pr}(F(D^{\prime})\in O)+\delta$ - where $\epsilon$ and $\delta$ are the privacy parameters that represent the privacy budget $\epsilon$ (measure of privacy loss), and the probability $\delta$ of the privacy being compromised, respectively [22]. The smaller the values of these parameters, the stronger the privacy guarantees. A DP algorithm $F$ is usually created out of an algorithm $F^{*}$ by adding noise proportional to the sensitivity of $F^{*}$, where sensitivity is maximum change in $F^{*}$’s output when comparing outputs on $D$ and $D^{\prime}$. The Gaussian Mechanism achieves this by adding noise drawn from a Gaussian (normal) distribution with 0 mean and with a standard deviation scaled to the $l_{2}$-sensitivity of $F^{*}$.
Synthetic Data Generation (SDG). Consider a database $D$ with $d$ features (attributes) denoted by $x=\{x_{1},x_{2},\ldots,x_{d}\}$. The domain for feature $x_{i}$ is a finite, discrete set represented by $\Omega_{i}$, i.e., $|\Omega_{i}|=\omega_{i}$. Let $\mathcal{Q}$ represent a collection of measurement sets where each set $q$ in $Q$ is a set of features to measure, i.e., $q\subseteq x$. A marginal on $q$ is denoted by $\mu_{q}(D)$ and simply refers to the number of occurrences in $D$ for each $t\in\Omega_{q}$ where $\Omega_{q}=\prod_{x_{i}\in q}\Omega_{i}$. Example: all 2-way marginals will consist of all $q$ where $|q|=2$, i.e., all 2-combinations of features such as $\{x_{i},x_{j}\}$ where $i,j\in\{1,\ldots,d\}$ and $i<j$. Informally, one can for instance think of the 2-way marginal on $\{\textit{Gender},\textit{Age}\}$ as a histogram over all possible combinations of gender and age values in $D$. “Select-measure-generate” SDG algorithms aim at creating synthetic tabular data with marginals that are as close as possible to the original data [23, 24, 25, 26]. In this paper, we use Private-PGM [27] as a prototypical algorithm from this family of SDGs. Private-PGM constructs undirected graphical models from DP noisy measurements over low-dimensional marginals, which facilitates the generation of new synthetic samples via sampling from the learned graphical model. It works on records with discrete attributes. Simply put, for each $q$, Private-PGM firstly computes DP marginals $\mu_{q}(D)+\mathcal{N}(0,\sigma^{2}_{q}I)$, where $\mathcal{N(.)}$ is Gaussian noise with scale $\sigma^{2}_{q}$ determined based on $(\epsilon,\delta)$. This is the measurement step. Then it estimates the joint marginal distribution that best explains all the noisy measurement. In parallel, it estimates the parameters of the graphical model using graph inference and learning algorithms such as belief propagation on a junction tree. This is the generate step. We refer to McKenna et al. [27] for more details.
@@ 3 Secure Generation and Publishing of Synthetic Data
2:Data custodians, $\mathbf{C}$, secret share their respective datasets with MPC servers, $\mathbf{S}$, as $D_{i}]\!]$ and thresholds of data quality $T_{i}]\!]$
[…]
@@
13:   $D_{train}]\!],[\![D_{test}]\!]\leftarrow$ getData([[D]][\![D]\!])
14:   $D_{train}]\!]\leftarrow\pi_{\mathsf{PREPROCESS}}([\![D_{train}]\!],(\epsilon_{p},\delta_{p}))$
15:   $\hat{D}_{train}]\!]\leftarrow\pi_{\mathsf{SDG}}([\![D_{train}]\!],(\epsilon_{s},\delta_{s}),\mathbf{H_{k}})$
16:   $r]\!]\leftarrow\pi_{\mathsf{EVAL}}([\![\hat{D}_{train}]\!],[\![D_{test}]\!],[\![D_{train}]\!])$
17:   $R[k]]\!]\leftarrow[\![r]\!]$
@@
19:  $M]\!]\leftarrow\pi_{\mathsf{AVG}}([\![R[k]]\!]|\forall k)$
@@
21:  for every data custodian $c$ get threshold $T_{c}]\!]$ do
22:   votes $\leftarrow$ votes $([\![M]\!<[\![T_{c}]\!]))$
@@
28:  ${D}]\!]\leftarrow\pi_{\mathsf{PREPROCESS}}([\![D]\!],(\epsilon_{p},\delta_{p}))$
29:  $\hat{D}]\!]\leftarrow\pi_{\mathsf{SDG}}([\![D]\!],(\epsilon_{s},\delta_{s}),\mathbf{H_{s}})$
[…]
@@
The data custodians $\mathbf{C}$ secret share their private data $D_{i}]\!]$ and the desired quality of synthetic data as thresholds $T_{i}]\!]$ to MPC servers $\mathbf{S}$ on Line 1, while servers set up based on the secret-shared inputs received from the data custodians and public inputs such as the dimensions of the data and the set of parameters to tune for. $\mathbf{S}$ then run $\pi_{\mathsf{CONCAT}}$ on Line 3 to concat the secret shares of the data from all $\mathbf{C}$ and hold the secret shares of [[D]][\![D]\!] at the end of this protocol, hence, mimicking the centralized setup, as if all the data was in one location.
@@
Parameter tuning is performed iteratively from Lines 6 – 25 until one of the following conditions is met: the hyperparameter set is exhausted, the maximum number of iterations is reached, or the desired quality of synthetic data is achieved. [^7] This can be easily modified to exhaustively search for all parameters, and then choose the best hyperparameter for which we obtained the best evaluation metrics. This would require large number of secure comparisons - comparing with thresholds and comparing with all the sets of averaged evaluation metrics. Each iteration begins with K-fold (Lines 8 – 17), creating train and test splits based on indices only with getData() on Line 12. $\mathbf{S}$ then run MPC protocol $\pi_{\mathsf{PREPROCESS}}$ to preprocess the secret shares of training data on Line 13 followed by an MPC protocol for $\pi_{\mathsf{SDG}}$ to train an SDG model on secret shares of preprocessed training data with the set of hyperparameter $\mathbf{H_{k}}$. On line 15, $\mathbf{S}$ run $\pi_{\mathsf{EVAL}}$ to compute secret shares of the pre-defined evaluation metrics [[r]][\![r]\!] based on secret shares of both the real $D_{train}]\!],[\!D_{test}]\!]$ and generated synthetic data $\hat{D}_{train}]\!]$. The value of [[r]][\![r]\!] is never published and hence need not be DP. Note that all of these steps perform computation on combined real data and hence are done in MPC. On line 18, $\mathbf{S}$ compute secret shares of K-fold evaluation [[M]][\![M]\!] using $\pi_{\mathsf{AVG}}$.
@@
Lines 19 – 23 determine if [[M]][\![M]\!] meets the thresholds of all the data custodians. To do so, assume that [[M]][\![M]\!] meets the threshold for all $\mathbf{C}$ on Line 19, i.e., votes $|\mathbf{C}|$. Line 21 decrements the votes if it does not meet the threshold of a data custodian, finally checking if the number of votes still equals $|\mathbf{C}|$ on Line 23. If yes, it is set to publish on Line 23, selecting the hyperparameters for this iteration as the final set of hyperparamters $\mathbf{H_{s}}$. The voting is done in a privacy-preserving way – (a) it does not require release of any evaluation metrics; and (b) the thresholds of each data custodian and the number of votes are kept confidential. Finally, on Line 27 – 29, the SDG pipeline is run on the combined data [[D]][\![D]\!], including reverse preprocessing on Line 29 to match the format of the real data that was secret shared initially. Line 30 publishes generated synthetic data of the same length as the combined data, which can be changed in the framework (e.g. on Line 28). Note that the framework is modular in nature and can be switched to suit any specific preprocessing or SDG algorithm or evaluation metrics [^8] The code for our framework is available at https://https://anonymous.4open.science/endsdg. Note that it is work in progress.. Once we have the MPC protocols in-place, one can easily modify to change the method of parameter search as long as it does not require access to private data.
@@ 5.1 Secure Preprocessing
1:Secret shares of one gene feature column of $D^{g}]\!]$ for gene $g$, the number of samples NN
2:Secret shares of one gene feature binned column $D^{g}]\!]$ for gene $g$, secret shares of mean $m^{g}\!]$ for gene $g$ for inverse discretization done later
@@
4:Create a copy of $D^{g}]\!]$ as $D^{g}_{copy}$}]\!]; Initialize [[Q]][\![Q]\!] of length 3
[…]
@@
[…]
@@
10:for $i\leftarrow 1$ to NN do $\triangleright$ Binning data into 4 bins with values 0,1,2,3
11:  $c_{0}]\!]\leftarrow\pi_{\mathsf{LT}}([\![D^{g}[i]]\!]$, [[Q[[\![Q[0]]])]]\!]); $c_{1}]\!]\leftarrow\pi_{\mathsf{LT}}([\![D^{g}[i]]\!]$, [[Q[[\![Q[1]]])]]\!]); $c_{2}]\!]\leftarrow\pi_{\mathsf{LT}}([\![D^{g}[i]]\!]$, [[Q[[\![Q[2]]])]]\!])
12:  $D^{g}[i]]\!]\leftarrow 3-[\![c_{0}]\!]-[\![c_{1}]\!]-[\![c_{2}]\!]$
@@
15:for $i\leftarrow 1$ to NN do $\triangleright$ Computing mean for inverse discretization
16:  for $b\leftarrow 0$ to 33 do
17:   $c]\!]\leftarrow\pi_{\mathsf{EQ}}([\![D^{g}[i]]\!],b)$
18:   $m^{g}[b]\!]\leftarrow[\![m^{g}[b]\!]+\pi_{\mathsf{MUL}}([\![c]\!],[\![D^{g}_{copy}[i]]\!]$)
19:   $^{g}[b]]\!]\leftarrow[\![ctr$$^{g}[b]]\!]+[\![c]\!]$
@@
Lines 13–22 compute the mean values for each bin index in a straightforward manner, using MPC protocols for equality testing $\pi_{\mathsf{EQ}}$ and multiplication $\pi_{\mathsf{MUL}}$ of secret-shared values (see Appendix A). Note that [[[\![ctr[$^{g}[b]]\!$] is used as a secret-shared accumulator variable for the number of values that fall into bin $b$.
Description of $\pi_{\mathsf{INV-BIN}}$. $\pi_{\mathsf{INV-BIN}}$ takes as input secret shares of binned [[D]][\![{D}]\!] and secret shares of all mean values $m^{g}]\!]$ and outputs the de-binned data. We replace the value $g_{i}$ with the mean following $\sum_{i=0}^{3}m^{g}[b]\cdot(g_{i}==b)$ on Lines 4–5. We note that line 4 can be further optimized by using the indicator polynomials described in Section 5.2.
@@
1:Secret shares of binned data of [[D]][\![D]\!], the number of samples NN, secret shares of mean $m^{g}\!]$ for all genes
2:Secret shares of de-binned data [[D]][\![D]\!]
3:for $i\leftarrow 1$ to NN do $\triangleright$ Binned Data into 4 bins with values 0,1,2,3
@@
5:   $x]\!]\leftarrow[\![D[i][g]]\!]$
6:   $c_{0}]\!]\leftarrow\pi_{\mathsf{EQ}}([\![x]\!],0)$; $c_{1}]\!]\leftarrow\pi_{\mathsf{EQ}}([\![x]\!],1)$; $c_{2}]\!]\leftarrow\pi_{\mathsf{EQ}}([\![x]\!],2)$; $c_{3}]\!]\leftarrow\pi_{\mathsf{EQ}}([\![x]\!],3)$
7:   $D[i][g]]\!]\leftarrow\pi_{\mathsf{MUL}}([\![c_{0}]\!],[\![m^{g}\!][0])+\pi_{\mathsf{MUL}}([\![c_{1}]\!],[\![m^{g}\!][1])+\pi_{\mathsf{MUL}}([\![c_{2}]\!],[\![m^{g}\!][2])+\pi_{\mathsf{MUL}}([\![c_{3}]\!],[\![m^{g}\!][3])$
@@ 5.2 Secure SDG
[…]
@@
1:Secret shares of binned data [[D]][\![D]\!], scale $\sigma$, number of patients NN, Domain of gene features $\Omega_{g}={0,1,2,3}$ and classes $\Omega_{y}={0,1,2,3,4}$
2:Secret shares of noisy marginals $\mu]\!]$ for 1-way and selected 2-way marginals
3:Initialize matrix $\mu_{g}]\!]$ of size $[d,4]$, array $\mu_{y}]\!]$ of size $[5]$ and matrix $\mu_{g,y}]\!]$ of size $[d,20]$
4:for $i\leftarrow 1$ to NN do
5:  $x]\!]\leftarrow[\![D[i][y]]\!] \triangleright$ Compute 1-way marginal for the label
6:  Initialize array [[L]][\![L]\!] of size 5
7:  $s_{0}]\!]\leftarrow[\![x]\!]$; $s_{1}]\!]\leftarrow[\![x]\!]-1$; $s_{2}]\!]\leftarrow[\![x]\!]-2$; $s_{3}]\!]\leftarrow[\![x]\!]-3$; $s_{4}]\!]\leftarrow[\![x]\!]-4$;
8:  $L[0]]\!]\leftarrow\pi_{\mathsf{MUL}}([\![s_{1}]\!],[\![s_{2}]\!],[\![s_{3}]\!],[\![s_{4}]\!],(1/24))$
9:  $L[1]]\!]\leftarrow\pi_{\mathsf{MUL}}([\![s_{0}]\!],[\![s_{1}]\!],[\![s_{3}]\!],[\![s_{4}]\!],(1/6))$
10:  $L[2]]\!]\leftarrow\pi_{\mathsf{MUL}}([\![s_{0}]\!],[\![s_{1}]\!],[\![s_{2}]\!],[\![s_{4}]\!],(1/2))$
11:  $L[3]]\!]\leftarrow\pi_{\mathsf{MUL}}([\![s_{0}]\!],[\![s_{1}]\!],[\![s_{2}]\!],[\![s_{4}]\!],(1/6))$
12:  $L[4]]\!]\leftarrow\pi_{\mathsf{MUL}}([\![s_{0}]\!],[\![s_{1}]\!],[\![s_{2}]\!],[\![s_{3}]\!],(1/24))$
@@
17:   $x]\!]\leftarrow[\![D[i][g]]\!]$
18:   Initialize array [[G]][\![G]\!] of size 5
19:   $s_{1}]\!]\leftarrow 1-[\![x]\!]$; $s_{11}]\!]\leftarrow[\![x]\!]-1$; $s_{2}]\!]\leftarrow 2-[\![x]\!]$; $s_{21}]\!]\leftarrow[\![x]\!]-2$; $s_{3}]\!]\leftarrow 3-[\![x]\!]$; $s_{31}]\!]\leftarrow[\![x]\!]-3$;
20:   $G[0]]\!]\leftarrow\pi_{\mathsf{MUL}}([\![s_{1}]\!],[\![s_{2}]\!],[\![s_{3}]\!],(1/6))$
21:   $G[1]]\!]\leftarrow\pi_{\mathsf{MUL}}([\![x]\!],[\![s_{2}]\!],[\![s_{3}]\!],(1/2))$
22:   $G[2]]\!]\leftarrow\pi_{\mathsf{MUL}}([\![x]\!],[\![s_{11}]\!],[\![s_{3}]\!],(1/2))$
23:   $G[3]]\!]\leftarrow\pi_{\mathsf{MUL}}([\![x]\!],[\![s_{11}]\!],[\![s_{22}]\!],(1/6))$
@@
35:for every value [[x]][\![x]\!] in $\mu_{g}]\!]$, $\mu_{l}]\!]$, $\mu_{l}g]\!]$ do $\triangleright$ Adding noise to guarantee DP
36:  $x]\!]\leftarrow[\![x]\!]+\sigma\cdot\pi_{\mathsf{GAUSS}}(0,1)$
@@
[…]